# Private Online Community Detection for Censored Block Models

### v1

@@ 5.1 Performance Analysis for the Proposed Method
In the proposed detection method, we have a sliding window estimate for the unknown community structures. We first analyze the sufficient conditions for exact community recovery under privacy constraints in Section 5.1.1.
5.1.1 Sufficient Condition for Exact Community Recovery
Theorem 5.2.
The graph perturbation based recovery algorithm via Warner’s randomized response satisfies $\epsilon$-edge CDP, $\forall\epsilon\geq\epsilon_{n}=\Omega(\log(n))$, $\forall a>\frac{2(n^{3/2}-n)}{(n-1)\log(n)}$, and it satisfies exact recovery if
@@ 5.1.1 Sufficient Condition for Exact Community Recovery
The proof is given in Appendix E.
Remark 5.3.
Observe that, if $\epsilon$ remains constant while $n$ increases, it follows that from (13), $\lim_{n\rightarrow\infty}\tilde{\zeta}=1/2$. This suggests that with a consistent $\epsilon$, the characteristics of both inter- and intra-community edges converge in the asymptotic limit, rendering exact recovery infeasible. However, Theorem 5.2 demonstrates that exact recovery can be achieved when the leakage grows logarithmically with $n$.
5.1.2 Necessary Condition for Exact Community Recovery
We subsequently introduce an information-theoretic lower bound that delineates the minimum necessary privacy levels for achieving exact community recovery in $\epsilon$-edge CDP frameworks under CBMs. This lower bound serves as a critical benchmark for assessing the effectiveness and limitations of privacy-preserving mechanisms in precise community detection within graph structures, particularly under the stringent conditions imposed by $\epsilon$-edge CDP.
Theorem 5.3 (Converse Result).
Suppose there exists an $\epsilon$-edge CDP mechanism such that for any labeling vector $\bm{\sigma}\in\{\pm 1\}^{n}$, on $\mathbf{A}\sim{\mathsf{CBM}}(\bm{\sigma},p,\zeta)$, output $\hat{\bm{\sigma}}\in\{\pm 1\}^{n}$ satisfying (3). Then, the privacy leakage must scale at least as
@@ 5.1.2 Necessary Condition for Exact Community Recovery
where $p^{\prime}=2p^{2}\zeta(\zeta-1)-(p-1)^{2}+1$, $p=a\log(n)/n$ and $\zeta\in(0,1/2)$.
The proof is given in Appendix F.
Remark 5.4.
@@
The lower bound derived in this context reveals the direct connection between the CBM parameters $a$ and $\zeta$, and the privacy budget $\epsilon$ required for exact community recovery. Distinctly, this lower bound decays slower than the bound derived from the traditional packing lower bound argument method, as discussed in [33]. In contrast to the constant nature of our derived lower bound, the packing argument suggests a more dynamic relationship, specifically, $\epsilon=\Omega\left({\log(n)}/{{n\choose 2}}\right)$, highlighting a less robust link with $n$. For smaller values of $\epsilon\leq 1$, the lower bound on the privacy budget in (20) scales as $\epsilon=\Omega\left({\log(n)}/{n}\right)$, which indicates a nuanced scaling behavior in comparison to the previously discussed bound. It is worth highlighting that the proposed graph perturbation mechanism requires $\epsilon=\Omega(\log(n))$. There remains an interesting opportunity to close this gap by devising more sophisticated mechanisms.
These results provide us sufficient and necessary conditions on the privacy level $\epsilon$ in order to guarantee efficient detection using privatized data.
5.1.3 False Alarm and Detection Delay
Under the exact recovery, we analyze the property of the estimated $\hat{\bm{\sigma}}$ and then the detection performance. All technical details can be found in Appendix G.
First of all, we have the following guarantee for the average run length that measures the false alarm performance.
Lemma 5.2 (Average Run Length).
@@ 5.1.3 False Alarm and Detection Delay
The proposed detection method (17) satisfies, for any detection threshold $b$,
${\mathbb{E}}_{\infty}[T]\geq e^{b}.$
It is obvious to see that we can choose the detection threshold $b=\log\gamma$ such that the average run length constraint ${\mathbb{E}}_{\infty}[T]\geq\gamma$ is satisfied. Therefore, we adopt this threshold in subsequent delay analysis, and provide an upper bound for the detection delay in the following.
Theorem 5.4 (WADD Upper Bound).
@@
Assume the exact recovery condition (19) is satisfied, we have the following upper bound on the worst-case performance of the detection procedure (17), when choosing $b=\log\gamma$, as $\gamma\to\infty$, and $w\tilde{I}_{0}=o(\log\gamma)$:
${\mathbb{E}}_{0}[T]\leq\frac{\log\gamma}{\tilde{I}_{0}}(1+o(1)),$ (21)
where $\tilde{I}_{0}$ is the KL divergence of the privatized post-change distribution from the privatized pre-change distribution defined in (18).
Remark 5.5.
We note that, compared with the results in [36], our setting and method here have two fundamental differences. First, the parameter to be estimated in our case, $\bm{\sigma}$, is discrete, while the maximum likelihood estimate in [36] was conducted for continuous values. Second, we use an approximate maximum likelihood estimator resulting from the semidefinite relaxations, whereas [36] employed the exact maximum likelihood estimator. These differences call for new analysis techniques to address discrete parameter spaces and the approximation error of semidefinite relaxations, which we provide in the appendix.
5.2 Delay Lower Bound with Private Constraint
@@
Our objective is to detect changes in dynamic networks under privacy constraints. The lower bound on the detection delay in Theorem 5.1 is the information-theoretic lower bound under the particular private mechanism used in Section 4.1. In this section, we further investigate the minimax lower bound for arbitrary private algorithms that satisfies $\epsilon$-edge LDP, as defined as
@@
where $\mathcal{Q}_{\epsilon}$ is the set of all privacy mechanisms that satisfy the $\epsilon$-edge LDP constraint, $\tau$ ranges over all stopping rules in the set $\mathcal{C}(\gamma)$ satisfying the ARL constraint; and the inner supremum denotes the worst-case detection delay over all possible change-point locations, with $\mathbb{E}_{\cdot,Q}$ denoting the expectation under the current privacy mechanism $Q$. We aim to provide a general lower bound to (22) and compare with the delay of our proposed method in the end.
To solve the above minimax lower bound, we first analyze the KL divergence after the privacy mechanism. We denote $p_{\infty}({\mathbf{A}})$ and $p_{0}({\mathbf{A}})$ as the probability distribution for the adjacency matrix ${\mathbf{A}}$ in the pre-change and post-change regime, respectively. And denote $\tilde{p}_{\infty}({\mathbf{A}})$ and $\tilde{p}_{0}({\mathbf{A}})$ as the corresponding two distributions for the privatized data distribution. Note that we have
${\mathsf{KL}}(\tilde{p}_{0}||\tilde{p}_{\infty})$ $=\mathbb{E}_{\tilde{p}_{0}}\bigg{[}\sum_{1\leq i<j\leq n}\log\frac{\tilde{p}_{0}(\tilde{A}_{ij})}{\tilde{p}_{\infty}(\tilde{A}_{ij})}\bigg{]}.$
Denote $q_{1}(z)=q(\tilde{A}_{ij}=z|A_{ij}=1)$ as the probability that the privatized edge equals to $z$ when the raw edge is $+1$, and similarly we define $q_{0}(z)$ and $q_{-1}(z)$. By the privacy constraint we have $\frac{q_{x}(z)}{q_{y}(z)}\in[e^{-\epsilon},e^{\epsilon}]$ for any $x,y\in\{-1,0,+1\}$. Based on this, and following [11], we have the following upper bound for the KL divergence after any $\epsilon$-edge LDP perturbation. Technical details can be found in Appendix H.
Lemma 5.3 (KL divergence after perturbation).
@@
${\mathsf{KL}}(\tilde{p}_{0}||\tilde{p}_{\infty})\leq\min\{4,e^{2\epsilon}\}(e^{\epsilon}-1)^{2}p^{2}(1-2\zeta)^{2}\bigg{[}{n\choose 2}-C_{\bm{\sigma}^{\text{pre}},\bm{\sigma}^{\text{post}}}\bigg{]}.$
The above Lemma implies that
@@
$\inf_{Q\in\mathcal{Q}_{\epsilon}}\inf_{\tau\in\mathcal{C}(\gamma)}\underset{\nu\geq 1}{\sup}\operatorname*{ess\,sup}\ \mathbb{E}_{\nu,Q}\left[(\tau-\nu+1)^{+}|\mathcal{F}_{\nu-1}\right]\geq\\ \frac{\log\gamma}{\min\{4,e^{2\epsilon}\}(e^{\epsilon}-1)^{2}p^{2}(1-2\zeta)^{2}({n\choose 2}-C_{\bm{\sigma}^{\text{pre}},\bm{\sigma}^{\text{post}}})}(1+o(1)).$ (23)
@@
Meanwhile, recall that in our privacy mechanism, we have that the KL divergence after the randomized perturbation becomes $\tilde{I}_{0}=\left(\log\frac{1+p(1-\zeta)(e^{\epsilon}-1)}{1+p\zeta(e^{\epsilon}-1)}\right)\cdot p(1-2\zeta)\frac{e^{\epsilon}-1}{e^{\epsilon}+2}({n\choose 2}-C_{\bm{\sigma}^{\text{pre}},\bm{\sigma}^{\text{post}}})\approx p(1-2\zeta)(e^{\epsilon}-1)\cdot p(1-2\zeta)\frac{e^{\epsilon}-1}{e^{\epsilon}+2}({n\choose 2}-C_{\bm{\sigma}^{\text{pre}},\bm{\sigma}^{\text{post}}})$ for small $\epsilon$ (i.e., strong privacy constraint), which matches with the denominator in (23) up to a constant factor.
6 Numerical Results

### v2

@@ IV-A Conditions for Exact Community Recovery
In the proposed detection method, we have an online estimate for the unknown community structure. We first analyze the sufficient conditions for exact community recovery under LDP privacy constraints.
Definition 4 (Exact Recovery).
$\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A})={\bm{\sigma}^{*}})\geq 1-o(1),$ (11)
where the probability is taken over both the randomness of the observation $\mathbf{A}$} and the estimator.
Theorem IV.1.
Suppose $\epsilon\geq\epsilon_{n}=\Omega(\log(n))$, $p=a\frac{\log(n)}{n}$, and $a>\frac{2(n^{3/2}-n)}{(n-1)\log(n)}$. The estimator from (3), using perturbed graph $\tilde{\mathbf{A}}$ under our Privacy Mechanism in Section III-A, satisfies exact recovery if
@@
The proof of this theorem is presented in Appendix C-A. Observe that, if $\epsilon$ remains constant while $n$ increases, it follows that from (6), $\lim_{n\rightarrow\infty}\tilde{\zeta}=1/2$. This suggests that with a consistent $\epsilon$, the characteristics of both inter- and intra-community edges converge in the asymptotic limit, rendering exact recovery infeasible. However, Theorem IV.1 demonstrates that exact recovery can be achieved when the leakage grows logarithmically with $n$. Our analysis also reveals an interesting trade-off between the privacy scaling coefficient $c$, where $\epsilon=c\log(n)$ and the CBM parameter $a$ as highlighted in Fig. 10 in the Appendix.
Sufficient Condition under Stability Mechanism
We next present a sufficient exact recovery condition under CDP for the perturbation stability-based recovery algorithm in the following theorem. We also plot and compare the threshold conditions for exact recovery under both mechanisms, as shown in Fig. 4.
Theorem IV.2.
The perturbation stability-based mechanism (described in Algorithm 2) satisfies $(\epsilon,\delta)$-edge CDP, and it satisfies exact recovery if
$a\big(\sqrt{1-\zeta}-\sqrt{\zeta}\big)^{2}>1,$ (13)
for $a>3/\epsilon$, $p=a\log(n)/n$, $\epsilon>0$, and $\delta=n^{-2}$.
Figure 4: Exact Recovery Regime that depends on $(a,\zeta)$, where $n=100$ and $\epsilon=\log(n)$.
We also present a sufficient condition for exact recovery for the computationally-efficient stability mechanism in the following theorem.
Theorem IV.3.
Algorithm 3 satisfies $(\epsilon,2\delta)$-edge CDP, and it satisfies exact recovery if
$a\big(\sqrt{1-\zeta}-\sqrt{\zeta}\big)^{2}>\max\left\{\frac{32\log(n)}{\epsilon},1\right\},$ (14)
for $p=a\log(n)/n$, $\epsilon>0$, and $\delta=n^{-2}$.
The proof follows along lines similar to those in our work in [34].
Remark IV.1 (Comparison with Graph Perturbation Mechanism).
A key advantage of the perturbation stability mechanism is its constant privacy leakage guarantee, i.e., $\epsilon=\mathcal{O}(1)$. However, it leads to a slightly higher running time and more strict recovery conditions. This highlights the fundamental tradeoffs between computational complexity, recovery conditions, and privacy guarantees.
Remark IV.2 (Memory Footprint and Scalability).
Efficient storage is critical when privacy mechanisms are applied to large graphs; the following bounds summarise the peak RAM usage of each step and how we scale beyond it.
  • Randomized‑response: A dense representation of the perturbed graph needs $\Theta(n^{2})$ bits, whereas a compressed edge list uses only $\Theta(m)$ words for sparse graphs.
  • Subsampling-stability mechanism: At any given time, only the current subsample $S\subseteq E$ of size $\lvert S\rvert=\beta\,m$ (for some $0<\beta<1$) is held in RAM. Hence, the peak memory cost is $\Theta\bigl(n+\beta\,m\bigr)$, which is linear in the graph size and strictly below the cost of storing the full edge list whenever $\beta<1$.
For massive graphs, we can further reduce resources by sketching [34] the graph to $\tilde{O}(n^{\alpha})$ nodes/edges with $\alpha<1$, running community detection on the sketch, and propagating the labels back. This lowers both time and memory by a factor of $n^{1-\alpha}$ with negligible loss in accuracy, and the same idea accelerates the SDP formulation even under the extra separation constraint induced by subsampling.
Necessary Condition for Exact Community Recovery
We subsequently introduce an information-theoretic lower bound that delineates the minimum necessary privacy levels for achieving exact community recovery in $\epsilon$-edge CDP frameworks. This lower bound serves as a critical benchmark for assessing the effectiveness and limitations of different privacy-preserving mechanisms.
Theorem IV.4 (Converse Result).
Suppose there exists an $\epsilon$-edge CDP mechanism such that for any labeling vector $\bm{\sigma}\in\{\pm 1\}^{n}$, on $\mathbf{A}\sim{\mathsf{CBM}}(\bm{\sigma},p,\zeta)$, output $\hat{\bm{\sigma}}\in\{\pm 1\}^{n}$ satisfying the exact recovery (11). Then, the privacy leakage must scale at least as
@@
where $p^{\prime}=2p^{2}\zeta(\zeta-1)-(p-1)^{2}+1$, $p=a\log(n)/n$, and $\zeta\in(0,1/2)$.
@@
The proof of this theorem is presented in Appendix C-B. The lower bound derived in this context reveals the direct connection between the CBM parameters $a$ and $\zeta$, and the privacy budget $\epsilon$ required for exact community recovery. Distinctly, this lower bound decays slower than the bound derived from the traditional packing lower bound argument method, as discussed in [37]. In contrast to the constant nature of our derived lower bound, the packing argument suggests a more dynamic relationship, specifically, $\epsilon=\Omega\left({\log(n)}/{{n\choose 2}}\right)$, highlighting a less robust link with $n$. For smaller values of $\epsilon\leq 1$, the lower bound on the privacy budget in (15) scales as $\epsilon=\Omega\left({\log(n)}/{n}\right)$, which indicates a nuanced scaling behavior in comparison to the previously discussed bound. It is worth highlighting that the proposed graph perturbation mechanism requires $\epsilon=\Omega(\log(n))$. There remains an interesting opportunity to close this gap by devising more sophisticated mechanisms.
IV-B Guarantees on Average Run Length and Detection Delay
In this subsection, we provide guarantees for the average run length and detection delay of both the procedure $T_{\rm L}$ under LDP setting in Equation (8) and $T_{\rm C}$ under CDP setting in Equation (10).
Lemma IV.1 (Average Run Length of $T_{\rm L}$).
@@ IV-B Guarantees on Average Run Length and Detection Delay
For a given detection threshold $b>0$, we have the test $T_{\rm L}$ in method (8) satisfies ${\mathbb{E}}_{\infty}[T_{\rm L}]\geq e^{b}$.
The proof of this lemma is presented in Appendix D-A. In online change detection problems, a pre-specified large constant $\gamma$ is typically imposed as a lower bound on the average run length to prevent frequent false alarms. Based on Lemma IV.1, we may choose the detection threshold $b=\log\gamma$ to guarantee ${\mathbb{E}}_{\infty}[T_{\rm L}]\geq\gamma$ is satisfied. We next present the theoretical guarantee of the detection effectiveness.
Theorem IV.5 (Detection Delay of $T_{\rm L}$).
@@
Assume the exact recovery condition (12) is satisfied, we have the following worst-case average detection delay of the detection procedure (8), when $b=\log\gamma$ and $\tilde{I}_{0}=o(\log\gamma)$ as $\gamma\to\infty$:
${\mathsf{WADD}}[T_{\rm L}]=\frac{\log\gamma}{\tilde{I}_{0}}(1+o(1)),$ (16)
where
$\tilde{I}_{0}=\frac{1}{2}(\log\frac{1-\tilde{\zeta}}{\tilde{\zeta}})\tilde{p}(1-2\tilde{\zeta})\left({n\choose 2}-\sum_{i<j}{\sigma}^{\text{pre}}_{i}{\sigma}^{\text{pre}}_{j}{\sigma}_{i}^{\text{post}}{\sigma}_{j}^{\text{post}}\right).$
[…]
Remark IV.3 (Comparison with Non-Private Detection).
As shown in Theorem D.1 in the Appendix D-C, the information-theoretic lower bound on the WADD is given by $\frac{\log\gamma}{I_{0}}(1+o(1))$, where $I_{0}$ denotes the KL divergence between the post- and pre-change data distributions [36]. This bound is known to be achieved by the classical CUSUM procedure and thus serves as the optimal performance benchmark for non-private detection. In contrast, the detection delay of our private method above is $\frac{\log\gamma}{\tilde{I}_{0}}(1+o(1))$, where $\tilde{I}_{0}<I_{0}$ due to the information loss induced by privacy-preserving graph perturbations. This highlights the inherent trade-off between privacy guarantees and detection efficiency. As the privacy requirement becomes weaker (i.e., larger $\epsilon$), $\tilde{I}_{0}$ approaches $I_{0}$, and the performance of our private detection method approaches that of the optimal non-private baseline.
Similar to the results above, we also present guarantees on the test $T_{\rm C}$ under CDP constraints.
Lemma IV.2 (Average Run Length of $T_{\rm C}$).
Assume $\epsilon>8\log\frac{1-\zeta}{\zeta}$. For a given detection threshold $b>0$, we have the test (10) satisfies ${\mathbb{E}}_{\infty}[T_{\rm C}]\geq\frac{1-(4C/\epsilon)^{2}}{1-(2C/\epsilon)^{2}}e^{b}$, where $C=2\log(({1-\zeta})/{\zeta})$.
The proof of this lemma is presented in Appendix D-B. Again, we may choose the threshold $b=\log\gamma+\log\frac{1-(2C/\epsilon)^{2}}{1-(4C/\epsilon)^{2}}$ to guarantee $\mathbb{E}_{\infty}[T_{\rm C}]\geq\gamma$ is satisfied. For a constant privacy leakage $\epsilon$, we have the resulted threshold $b=\log\gamma\cdot(1+o(1))$.
Theorem IV.6 (Detection Delay of $T_{\rm C}$).
Assume the exact recovery condition (14) is satisfied and $\epsilon>8\log\frac{1-\zeta}{\zeta}$, we have the following worst-case average detection delay of the detection procedure (10), when $b=\log\gamma+\log\frac{1-(2C/\epsilon)^{2}}{1-(4C/\epsilon)^{2}}$ and $I_{0}=o(\log\gamma)$ as $\gamma\to\infty$:
${\mathsf{WADD}}[T_{\rm C}]=\frac{\log\gamma}{I_{0}}(1+o(1)),$ (17)
where
${I}_{0}=\frac{1}{2}(\log\frac{1-\zeta}{\zeta})p(1-2\zeta)({n\choose 2}-\sum_{i<j}{\sigma}^{\text{pre}}_{i}{\sigma}^{\text{pre}}_{j}{\sigma}_{i}^{\text{post}}{\sigma}_{j}^{\text{post}}).$
The proof of this theorem is presented in Appendix D-D. We note that ${I}_{0}$ is the KL divergence of the post- and pre-change $\mathsf{CBM}$} distributions without perturbation with ${I}_{0}>\tilde{I}_{0}$. Consequently, the delay in (17) is smaller than that in (16) (and $\tilde{I}_{0}\approx I_{0}$ for $\epsilon$ sufficiently large). This is consistent with the intuition that CDP is less stringent than LDP constraints, thus having less impact on the degradation of detection performance. On the other hand, the detection procedure under CDP setting requires more strict recovery conditions and a slightly higher running time caused by computation of stability.
IV-C Delay Lower Bound with Private Constraint
IV-C1 Delay Lower Bound under $\epsilon$-edge LDP
@@
In this section, we investigate the minimax lower bound to the detection delay for arbitrary private algorithms that satisfies $\epsilon$-edge LDP,
@@
where $\mathcal{Q}_{\epsilon}$ is the set of all privacy mechanisms that satisfy the $\epsilon$-edge LDP constraint, $\tau$ ranges over all stopping rules in the set $\mathcal{C}(\gamma)$ satisfying the ARL constraint $\mathbb{E}_{\infty}[\tau]\geq\gamma$; and the inner supremum denotes the worst-case detection delay over all possible change-point locations, with $\mathbb{E}_{\cdot,Q}$ denoting the expectation under the privacy mechanism $Q$. We aim to provide insights regarding a general lower bound to (18) and compare with the delay of our proposed method in the end.
To solve the above minimax lower bound, we first analyze the KL divergence after the privacy mechanism. We denote $p_{\infty}({\mathbf{A}})$ and $p_{0}({\mathbf{A}})$ as the probability distribution for the adjacency matrix $\mathbf{A}$} in the pre-change and post-change regime, respectively. And denote $\tilde{p}_{\infty}({\mathbf{A}})$ and $\tilde{p}_{0}({\mathbf{A}})$ as the corresponding two distributions for the privatized data distribution. Following [13], we have the following upper bound for the KL divergence after any $\epsilon$-edge LDP perturbation.
Lemma IV.3 (KL divergence after perturbation).
@@
${\mathsf{KL}}(\tilde{p}_{0}||\tilde{p}_{\infty})\leq C_{\epsilon}(e^{\epsilon}-1)^{2}p^{2}(1-2\zeta)^{2}\bigg[{n\choose 2}-C_{\bm{\sigma}^{\text{pre}},\bm{\sigma}^{\text{post}}}\bigg],$
with $C_{\epsilon}=\min\{4,e^{2\epsilon}\}$, $C_{\bm{\sigma}^{\text{pre}},\bm{\sigma}^{\text{post}}}:=\sum_{i<j}{\sigma}^{\text{pre}}_{i}{\sigma}^{\text{pre}}_{j}{\sigma}_{i}^{\text{post}}{\sigma}_{j}^{\text{post}}$.
The proof of this theorem is presented in Appendix E-A. Since the optimal worst-case average detection delay, under ARL constraint $\gamma$, for detecting distribution shifts from $\tilde{p}_{\infty}$ to $\tilde{p}_{0}$ is known to be $({\log\gamma})/{{\mathsf{KL}}(\tilde{p}_{0}||\tilde{p}_{\infty})}\cdot(1+o(1))$ [36], the above Lemma implies that
@@
$\inf_{Q\in\mathcal{Q}_{\epsilon}}\inf_{\tau\in\mathcal{C}(\gamma)}\underset{\nu\geq 1}{\sup}\operatorname*{ess\,sup}\ \mathbb{E}_{\nu,Q}\left[(\tau-\nu+1)^{+}|\mathcal{F}_{\nu-1}\right]\geq\\ \frac{\log\gamma}{C_{\epsilon}(e^{\epsilon}-1)^{2}p^{2}(1-2\zeta)^{2}({n\choose 2}-C_{\bm{\sigma}^{\text{pre}},\bm{\sigma}^{\text{post}}})}(1+o(1)).$ (19)
@@
Meanwhile, recall that in our privacy mechanism, we have that the KL divergence after the randomized perturbation becomes $\tilde{I}_{0}=\left(\log\frac{1+p(1-\zeta)(e^{\epsilon}-1)}{1+p\zeta(e^{\epsilon}-1)}\right)\cdot p(1-2\zeta)\frac{e^{\epsilon}-1}{e^{\epsilon}+2}({n\choose 2}-C_{\bm{\sigma}^{\text{pre}},\bm{\sigma}^{\text{post}}})\approx p(1-2\zeta)(e^{\epsilon}-1)\cdot p(1-2\zeta)\frac{e^{\epsilon}-1}{e^{\epsilon}+2}({n\choose 2}-C_{\bm{\sigma}^{\text{pre}},\bm{\sigma}^{\text{post}}})$ for small $\epsilon$, which matches with the denominator in (19) up to a constant factor. However, for large $\epsilon$ values, there still remains an opportunity to close the gap between the delay of our proposed algorithm and the lower bound in (19) by developing more efficient private perturbation mechanisms.
IV-C2 Delay Lower Bound under $(\epsilon,\delta)$-edge CDP
We analyze the lower bound on detection delay under CDP settings, focusing on a specific scenario where the detection problem is performed based on the output of binary hypothesis tests on individual observations at each time. More specifically, we consider a hypothesis test $\mathcal{T}$ with binary output $\{0,1\}$, where $\mathcal{T}({\mathbf{A}})=1$ implies there is a change (i.e., accept the alternative hypothesis $H_{1}$ that the change has happened) and vice versa, based on the observed graph $\mathbf{A}$}. Now, for any two adjacent graphs ${\mathbf{A}},{\mathbf{A}}^{\prime}$, an $(\epsilon,\delta)$-edge CDP test $\mathcal{T}$ should satisfies
${\operatorname{Pr}(\mathcal{T}(\mathbf{A}^{\prime})=a)}\leq e^{\epsilon}{\operatorname{Pr}(\mathcal{T}({\mathbf{A}})=a)}+\delta,\text{ for }a=0,1.$
Then we have the following Lemma that quantifies the performance of the $(\epsilon,\delta)$-edge CDP test $\mathcal{T}$.
Lemma IV.4.
For an $(\epsilon,\delta)$-edge CDP binary hypothesis test $\mathcal{T}$, we have
${\operatorname{Pr}(\mathcal{T}({\mathbf{A}})=1|p_{0})}-{\operatorname{Pr}(\mathcal{T}({\mathbf{A}})=1|p_{\infty})}$
$\leq\frac{1}{\sqrt{2}}\times\left(\frac{e^{R\epsilon}-1}{e^{R\epsilon}+1}\right)\times\left(1+\frac{2\delta}{e^{\epsilon}-1}\right)\times\sqrt{{\mathsf{KL}}(p_{0}||p_{\infty})},$ (20)
where $p_{\infty}({\mathbf{A}})$ and $p_{0}({\mathbf{A}})$ denote the probability distribution for the adjacency matrix $\mathbf{A}$} in the pre-change and post-change regime, respectively, and $R=2^{{n\choose 2}}$.
The proof of this lemma is presented in Appendix E-B.
We note that for non-private binary hypothesis test, we have the following bound
${\operatorname{Pr}(\mathcal{T}({\mathbf{A}})=1|p_{0})}-{\operatorname{Pr}(\mathcal{T}({\mathbf{A}})=1|p_{\infty})}\leq{\mathsf{TV}}(p_{0}||p_{\infty}).$
Indeed, the summation of type-I and type-II errors of any test is lower bounded by $1-{\mathsf{TV}}(p_{0}||p_{\infty})$, with this minimum value obtained by the Neyman-Pearson test. Compared with the above Equation (E-B), we observe that the additional constant $\frac{1}{\sqrt{2}}\frac{e^{R\epsilon}-1}{e^{R\epsilon}+1}\left(1+\frac{2\delta}{e^{\epsilon}-1}\right)$ can be viewed as the performance loss due to the privacy constraints.
We then present the general analysis of the information bound on the detection delay. Based on the converse result above for private hypothesis testing for a single observation, we consider the scenario that at each time $k$, an $(\epsilon,\delta)$-edge CDP test $\mathcal{T}$ is performed and the binary output is $O_{k}=\mathcal{T}({\mathbf{A}}_{k})\in\{0,1\}$. The detection is performed via the sequence of binary outputs $\{O_{k},k\in\mathbb{N}\}$.
Proposition IV.1.
When using the sequence of binary outputs $\{O_{k},k\in\mathds{N}\}$ for detection, with each $O_{k}$ resulted from an $(\epsilon,\delta)$-edge CDP hypothesis test $\mathcal{T}$ applied on sample ${\mathbf{A}}_{k}$, the detection delay is lower bound by,
${\mathsf{WADD}}\geq\frac{\log\gamma}{\frac{1}{\alpha_{0}}\left(\frac{e^{R\epsilon}-1}{e^{R\epsilon}+1}\right)^{2}\left(1+\frac{2\delta}{e^{\epsilon}-1}\right)^{2}{\mathsf{KL}}(p_{0}||p_{\infty})}(1+o(1)),$
as $\gamma\to\infty$, where $\alpha_{0}:=\operatorname{Pr}(\mathcal{T}({\mathbf{A}})=1|p_{\infty})>0$.
We prove this proposition in Appendix E-C. Recall that the information-theoretic lower bound to the asymptotic delay in the non-private setting is $(\log\gamma)/{\mathsf{KL}}(p_{0}||p_{\infty})$ [36]. Compared with the delay lower bound above, we note that the privacy impact is reflected in the constant term in the denominator that depends on $\epsilon$ and $\delta$.
IV-D Converse: How Many Graphs are Needed for Detection with Privacy Constraint?
In this subsection, we present a necessary condition for the number of graphs required to achieve the exact recovery condition under $\epsilon$-edge DP for arbitrary $\epsilon$. It is worth highlighting that the community recovery problem becomes easier as more graphs are observed and utilized together within the recovery algorithm.
Theorem IV.7.
For the CBM model, the number of observed graphs needed for exact recovery while satisfying $\epsilon$-edge DP should be at least
$w\geq\max\left\{\tilde{c}_{1}\times\frac{e^{\epsilon}}{e^{\epsilon}-1},\tilde{c}_{2}\times\log(n)\right\},$ (21)
for an arbitrary $\epsilon>0$, $\tilde{c}_{1}=1-n^{-\Omega(1)}$ and $\tilde{c}_{2}\triangleq{(4-4n^{-\Omega(1)})}/{\left(1-2-2n^{-\Omega(1)}\right)^{2}}$.
The proof of this theorem is presented in Appendix E-D. Here, the first term in the argument of (21) represents the edge privacy impact on the required window size, while the second term reflects the statistical limit required for exact recovery and follows from applying the basic Chernoff-Hoeffding concentration inequality. The above condition shows that the privacy impact is dominant in the high-privacy regime, where $\epsilon$ is small. Conversely, in the low-privacy regime, the lower bound on $w$ is mainly governed by the fundamental statistical limit, i.e., $w=\Omega(\log(n))$. This dual dependency ensures that the window size $w$ adapts appropriately to balance both privacy and recovery accuracy requirements.
The above theorem demonstrates that exact recovery is possible for arbitrary $\epsilon>0$ when $w>1$, unlike the previous case where $w=1$ (i.e., a single observed graph) in Theorem IV.4. Consequently, in order to achieve exact recovery, we may use the most recent $w$ graphs to estimate the unknown community and then calculate the detection statistics. We would like to highlight that while using additional graphs can be beneficial for improving recovery, it may be costly to store the most recent $w$ observations in online change detection especially when the required window size $w$ is large.
V Numerical Results
Refer to caption Refer to caption Refer to caption Refer to caption
Case 1 Case 2
Figure 5: Sample statistics trajectories and detection delay of the method $T_{\rm L}$ in (8) for both cases.
Refer to caption Refer to caption Refer to caption Refer to caption
Case 1 Case 2
Figure 6: Sample statistics trajectories and detection delay of the method $T_{\rm C}$ in (10) for both Cases.